OpenStudy (anonymous):

Help? I will fan and medal. Step by step so I understand :D 1. Put the following equation in slope-intercept form: 3x + 12y = -4 2. Put the following equation in standard form: 6y = -3x + 12

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

@jim_thompson5910 ??

jimthompson5910 (jim_thompson5910):

they want you to isolate y in `3x + 12y = -4`. Do you know how to do that?

OpenStudy (anonymous):

Not really good at it :? math is my worst subject

jimthompson5910 (jim_thompson5910):

ok the first thing you do is subtract 3x from both sides 3x + 12y = -4 3x + 12y-3x = -4-3x 12y = -3x-4 make sense?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

then you divide both sides by 12 to fully get y all by itself 12y = -3x-4 12y/12 = (-3x-4)/12 y = (-3x-4)/12 y = (-3x)/12 - 4/12 y = (-1/4)x - 1/3

jimthompson5910 (jim_thompson5910):

y = (-1/4)x - 1/3 is in y = mx+b form m = -1/4 is the slope b = -1/3 is the y intercept

jimthompson5910 (jim_thompson5910):

y = mx+b is slope intercept form

OpenStudy (anonymous):

Thanks :D

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

So do the same to the other one?

jimthompson5910 (jim_thompson5910):

IF they wanted it in slope intercept form, then yeah you would follow the same steps

jimthompson5910 (jim_thompson5910):

but they don't want it in slope intercept form, they want it in standard form

jimthompson5910 (jim_thompson5910):

Standard form is Ax + By = C Example: 2x + 3y = 10

OpenStudy (anonymous):

Oh Ok I think I got this part! Thanks so much!!

jimthompson5910 (jim_thompson5910):

you're welcome

jimthompson5910 (jim_thompson5910):

post what you get so I can check